OpenStudy (anonymous):

Use the quadratic formula to solve 2y2 – 10y + 8 = 0. A. {–4, –1} B. {4, –1} C. {–4, 1} D. {4, 1}

OpenStudy (jadeishere):

\[x = \frac{ 10 \pm \sqrt{100 - 4(2)(8)} }{ 4}\] \[x = \frac{ 10 \pm \sqrt{36} }{ 4 }\] \[x = \frac{ 10 - 6 }{ 4 } x = \frac{ 10 + 6 }{ 4 }\] \[x =\frac{ 4 }{ 4 } or x = \frac{ 16 }{ 4 }\] so x = 4 and x = 1

OpenStudy (anonymous):

Use the quadratic formula to solve 2y2 + 3y = 2. A. {–½, 2} B. {–½, –2} C. {½, 2} D. {½, –2}

OpenStudy (jadeishere):

\[x = \frac{ -3 \pm \sqrt{9 - 4(2)(-2)} }{ 4 }\]

OpenStudy (jadeishere):

\[x = \frac{ -3 \pm \sqrt{9 + 16} }{ 4}\]

OpenStudy (jadeishere):

\[x = \frac{ -3 \pm \sqrt{25}}{ 4}\] \[x = \frac{ -3 + 5 }{ 4 } or x = \frac{ -3 - 5 }{4 }\] \[x = \frac{ 2 }{ 4 } or x = \frac{ -8 }{ 4 }\]

OpenStudy (jadeishere):

given that, x = 1/2 and x = -2

OpenStudy (anonymous):

Use the quadratic formula to solve x2 + 7x + 8 = 0. Estimate irrational solutions to the nearest tenth. A. {–12, 5} B. {–11.1, –2.9} C. {–5.6, –1.4} D. {–3.9, 12.1}

OpenStudy (jadeishere):

just plug in the values to the formula, I've already done wrong by not properly explaining what I did